Based on the following data, develop a forecast for the expected demand for Year 2. Assume...

  1. Based on the following data, develop a forecast for the expected demand for Year 2. Assume the demand for year two will be 86,000 units. Please round to two decimal points in your answers. This is a seasonal demand forecast.

Year 1

Past Sales (Units)

Quarter 1

20,000

Quarter 2

26,000

Quarter 3

18,400

Quarter 4

19,800

Answer (You may illustrate this in tables if you wish, but show your calculation within the table)
Answer #1

Based on the data given,

Total Demand in Year 1 is (20000 + 26000 + 18400 + 19800)= 84,200

Average Demand per quarter in Year 1 is 84200/ 4 = 21050

Hence we will calculate the seasonal factor.

Year1 Sales Seasonal Factor
Q1 20000 20000/21050=0.9501
Q2 26000 26000/21050=1.2352
Q3 18400 18400/21050=0.8741
Q4 19800 19800/21050=0.9406

Now as given total demand for Year 2 is 86000

Average demand for Year 2 is 86000/4=21500

Basis this we calculate the forecast for Year 2.

Forecast for a quarter = Seasonal factor for the quarter * average demand for year 2

Year2 Forecast
Q1 0.9501*21500=20427.55
Q2 1.2352*21500=26555.82
Q3 0.8741*21500=18793.35
Q4 0.9406*21500=20223.28

Please give thumbs up if you like my answer. Thanks!